# TensoSDF: Roughness-aware Tensorial Representation for Robust Geometry and Material Reconstruction

### v1

@@ 6.3. Results on real data
Figure 5. Ablation studies on the geometry reconstruction. Introducing the TensoSDF (Figure (b)) can reconstruct more details than NeRO (Liu et al., 2023) (Figure (a)) but with problematic surfaces. After combining the radiance field and reflectance field with a fixed balancing weight of 0.5 (Figure (c)), the surface quality can be improved, but still with some flaws. Finally, using the roughness as the weight (Figure (d)) achieves the best quality.
6.4. Ablation study
We validate the key components of our method in Fig. 5 step by step. Starting from NeRO (Liu et al., 2023), we replace its pure MLP-based geometry representation with our TensoSDF. The TensoSDF can reconstruct more geometric details and significantly improves the quality (1.65 in MAE). Moreover, TensoSDF reduces the training time to about 50%, from 8h to less than 4h for each scene in our synthetic datasets. However, the reconstructed geometry still exhibits a noticeable dent on the surface. Then, after introducing both radiance and reflectance fields with a constant weight (0.5 in practice), the surface artifacts are alleviated. Finally, replacing the constant weight with the estimated roughness forms our complete solution and closely matches the ground truth.
6.5. Discussion and limitations
Although the high-capacity TensoSDF improves the geometric details and speeds up the training, it still introduces some limitations. One problem is the trade-off between fine details and high-frequency noise. Explicit representations are more prone to over-fitting locally, which inevitably leads to noise. A common mitigation is to use additional smooth regularization, which also causes the loss of details. It is worth to analyze this trade-off theoretically. The other problem is the storage cost. Even though the tensorial representation is compact compared to volume-based representations, the storage cost is higher than MLP-based methods. This problem can be alleviated by some implementation strategies, e.g., pruning. It would be interesting to explore more compact and powerful representations.
@@ 7. Conclusion and Future work
There are still many potential future researching directions. Regarding the training time, combining our method with the popular 3D Gaussian splatting (Kerbl et al., 2023) may reduce the training time further. For the reconstruction quality, the efficient tensorial representation provides an opportunity to jointly update the geometry and material in the second stage, which may improve the reconstruction quality. From the perspective of diversity, reconstructing translucent objects is also worth researching since there are many translucent objects in reality, such as jade, wax figures, etc. However, translucent materials need to consider higher dimensions, and designing an efficient formulation is the key.

### v2

@@ 6.5. Discussion and limitations
Size (MB)$\downarrow$ 8.3 70.5 17.9 37.6 109.0
@@ 7. Conclusion and Future work
There are still many potential future researching directions. Regarding the training time, combining our method with the popular 3D Gaussian splatting (Kerbl et al., 2023) may reduce the training time further. For the reconstruction quality, the efficient tensorial representation provides an opportunity to jointly update the geometry and material in the second stage, which may improve the reconstruction quality. From the perspective of diversity, reconstructing translucent objects is also worth researching. However, translucent materials need to consider higher dimensions, and designing an efficient formulation is the key. Our method currently focuses on single-object reconstruction with multi-view inputs. It would be interesting to apply our method on multi-object environments or few-shot reconstruction.
Acknowledgements.
We thank the reviewers for the valuable comments, and XiaoLong Wu for his help in our video. This work has been partially supported by the National Science and Technology Major Project under grant No. 2022ZD0116305 and National Natural Science Foundation of China under grant No. 62272275 and 62172220.